Constants PartA In an internal combustion engine, air at atmospheric pressure and a temperature of about 20 °C is compressed in the cylinder by a piston to 5.0 Estimate the temperature of the compressed air, assuming the pressure reaches 38 atm of its original volume (compression ratio 8.0). oC Submit est Ans Provide Feedback
Answer #1

Solution :-

Given in the question,

Temperature(T1) = 20°c +273.15 = 293.15

P2= 38atm

V 2= 1/8

P1V1/T1 = P2V2/T2

=> 1 x 1/(293.15) = 38 x (1/8)/T2

=> T2 = (38/8) x 293.15 = 1392.4625 K = 1392.4625 - 273.15 = 1119.3125 C